# Efficient designs for multivariate crossover trials

### v1

@@ 3.2 Generalized markov-type covariance structure
$\bm{{C}}_{d^{*}(s2)}$ $=\frac{n}{t-1}\begin{bmatrix}\bm{\varLambda}_{1}&\bm{\varLambda}_{2}\\ \bm{\varLambda}_{2}&\bm{\varLambda}_{4}\end{bmatrix},$ (3.7)
where the matrices $\bm{\varLambda}_{1}$, $\bm{\varLambda}_{2}$ and $\bm{\varLambda}_{4}$ are given as
$\begin{split}\bm{\varLambda}_{1}&=\left(tr\left(\bm{V}_{1}^{*}\right)+\frac{\bar{\rho}^{2}}{\sigma_{12}}tr\left(\bm{V}_{R}^{*}\right)-\frac{\left(tr\left(\bm{V}_{1}^{*}\bm{\psi}\right)\right)^{2}}{tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)}-\frac{\bar{\rho}^{2}}{\sigma_{12}}\frac{\left(tr\left(\bm{V}_{R}^{*}\bm{\psi}\right)\right)^{2}}{tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)}\right)\bm{H}_{t},\\ \bm{\varLambda}_{2}&=-\frac{\bar{\rho}}{\sigma_{12}}\left(tr\left(\bm{V}_{R}^{*}\right)-\frac{\left(tr\left(\bm{V}_{R}^{*}\bm{\psi}\right)\right)^{2}}{tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)}\right)\bm{H}_{t},\text{ and}\\ \bm{\varLambda}_{4}&=\frac{1}{\sigma_{12}}\left(tr\left(\bm{V}_{R}^{*}\right)-\frac{\left(tr\left(\bm{V}_{R}^{*}\bm{\psi}\right)\right)^{2}}{tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)}\right)\bm{H}_{t}.\end{split}$ (3.8)
@@
Since both conditions are not satisfied (see proof of Remark 3.2), we instead focus on efficient designs. In order to check whether $d^{*}$ is an efficient design, we obtain an upper bound of $tr\left(\bm{C}_{d(s2)}\right)$. If $d^{*}$ attains this upper bound then we say that $d^{*}$ is an efficient design. However, obtaining an upper bound for $tr\left(\bm{C}_{d(s2)}\right)$ is computationally challenging, so we instead consider $\bm{\tilde{C}}_{d(s2)}$ (see Lemma S.5), the information matrix for the direct effects with no period effects, and show that $\bm{{C}}_{d(s2)}\leq\bm{\tilde{C}}_{d(s2)}$ with equality if $d$ is uniform on periods (see Lemmas S.5 and S.7).
@@
$\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}=\{d:d\in\varOmega_{t,n=\lambda t(t-1),p=t},\text{ }t\geq 3\text{ and }d\text{ is a binary design}\}.$
Also, for $p=t$, a binary design is a design uniform on subjects, and thus $\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$ is indeed a class of designs with $p=t$ which are uniform on subjects.
@@
$\left(tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)+\frac{1+\bar{\rho}^{2}}{\sigma_{12}}tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)\right)^{-1}\Bigg{)}$. Here, $\bm{V}_{1}^{*}$, $\bm{V}_{R}^{*}$, $\bar{\rho}$ and $\sigma_{12}$ are as defined before.
@@
$\begin{split}tr\left(\bm{\tilde{C}}_{d(s2)(11)}\right)&=tr\left(\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\varOmega}_{1}\right)\bm{T}_{d}\right)+tr\left(\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\varOmega}_{4}\right)\bm{T}_{d}\right),\\ tr\left(\bm{\tilde{C}}_{d(s2)(12)}\right)&=tr\left(\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\varOmega}_{1}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\right)+tr\left(\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\varOmega}_{4}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\right),\text{ and}\\ tr\left(\bm{\tilde{C}}_{d(s2)(22)}\right)&=tr\left(\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{1}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\right)+tr\left(\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{4}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\right),\end{split}$ (3.13)
from Lemma S.5. Using Remark S.1, it is observed that $\bm{{1}}_{p}^{{}^{\prime}}\bm{\varOmega}_{1}=\bm{{1}}_{p}^{{}^{\prime}}\bm{\varOmega}_{2}=\bm{{1}}_{p}^{{}^{\prime}}\bm{\varOmega}_{4}=\bm{{0}}_{1\times p}$ and
$\begin{split}tr\left(\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\varOmega}_{1}\right)\bm{T}_{d}\right)&=n\times tr\left(\bm{\varOmega}_{1}\right),\\ tr\left(\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\varOmega}_{4}\right)\bm{T}_{d}\right)&=n\times tr\left(\bm{\varOmega}_{4}\right),\\ tr\left(\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\varOmega}_{1}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\right)&=n\times tr\left(\bm{\varOmega}_{1}\bm{\psi}\right),\\ tr\left(\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\varOmega}_{4}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\right)&=n\times tr\left(\bm{\varOmega}_{4}\bm{\psi}\right),\\ tr\left(\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{1}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\right)&=n\times tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{1}\bm{\psi}\right),\text{ and}\\ tr\left(\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{4}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\right)&=n\times tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{4}\bm{\psi}\right).\end{split}$ (3.14)
@@
$c_{22(22)}\big{)}$, where $c_{11(11)}=tr\left(\bm{\varOmega}_{1}\right)$, $c_{11(22)}=tr\left(\bm{\varOmega}_{4}\right)$, $c_{12(11)}=tr\left(\bm{\varOmega}_{1}\bm{\psi}\right)$, $c_{12(22)}=tr\left(\bm{\varOmega}_{4}\bm{\psi}\right)$, $c_{22(11)}=tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{1}\bm{\psi}\right)$ and $c_{22(22)}=tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{4}\bm{\psi}\right)$. From Lemma S.4, $\bm{\varOmega}_{1}=\bm{V}_{1}^{*}+\frac{\bar{\rho}^{2}}{\sigma_{12}}\bm{V}_{R}^{*}$ and $\bm{\varOmega}_{4}=\frac{1}{\sigma_{12}}\bm{V}_{R}^{*}$, where $\bar{\rho}=\rho\sqrt{\frac{\sigma_{22}}{\sigma_{11}}}$ and $\sigma_{12}=\sigma_{22}(1-\rho^{2})$. Using these expressions in $c_{22(11)}$ and $c_{22(22)}$, we obtain $c_{22(11)}=tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)+\frac{\bar{\rho}^{2}}{\sigma_{12}}tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)$ and $c_{22(22)}=\frac{1}{\sigma_{12}}tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)$. Using Lemma S.8, $|\bar{\rho}|>0$ and $\sigma_{12}>0$ gives $c_{22(11)}$, $c_{22(22)}>0$, implying $c_{22}=n\left(c_{22(11)}+c_{22(22)}\right)>0$. Combining (3.11), (3.12), (3.13) and (3.14), we obtain
@@
Note that since $c_{22}>0$, $c_{22}^{+}$ has been replaced by $c_{22}^{-1}$. Using the expression of $c_{11}$, $c_{12}$ and $c_{22}$, and the expressions of $\bm{\varOmega}_{1}$ and $\bm{\varOmega}_{4}$ in (3.15), we prove (3.10). ∎
@@
Let $d^{*}\in\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$ be a design given by $OA_{I}\left(n=\lambda t\left(t-1\right),p=t,t,2\right)$, where $\lambda$ is a positive integer and $t\geq 3$. Then $d^{*}$ attains $u\left(t,n,p,\sigma_{11},\sigma_{22},\rho,\bm{V}_{1},\bm{V}_{R}\right)$ as given in Lemma 3.1, where $n=\lambda t\left(t-1\right)$ and $p=t$, if and only if $tr\left(\bm{V}_{1}^{*}\bm{\psi}\right)tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)=tr\left(\bm{V}_{R}^{*}\bm{\psi}\right)tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)$. Here, $\bm{V}_{1}^{*}$ and $\bm{V}_{R}^{*}$ are defined in Lemma 3.1.
@@
The proof of the above remark is given in Appendix B. ∎
@@ 3.2.1 Study of design efficiency for the markovian covariance structure
In this section, for various choices for $\bm{V}_{1}$ and $\bm{V}_{R}$, we evaluate the performance of $d^{*}$ on the basis of the upper bound of $tr\left(\bm{C}_{d(s2)}\right)$. For the purpose of simplicity, we denote the upper bound given in Lemma 3.1 by $u$. Different combinations of $\bm{V}_{1}$ and $\bm{V}_{R}$ are considered:
Case 1:
$\bm{V}_{1}$ is $AR(1)$ and $\bm{V}_{R}=\bm{I}_{p}$.
Case 2:
$\bm{V}_{1}$ is tridiagonal and $\bm{V}_{R}=\bm{I}_{p}$.
Case 3:
$\bm{V}_{1}=\sigma_{11}\bm{I}_{p}$, where $\sigma_{11}>0$, and $\bm{V}_{R}$ is $AR(1)$.
Case 4:
$\bm{V}_{1}=\sigma_{11}\bm{I}_{p}$, where $\sigma_{11}>0$, and $\bm{V}_{R}$ is tridiagonal.
For our illustrations, $p=t\in\{3,4\}$ is used. For application of Lemma 3.1, we are required to find $\bm{V}_{1}^{*}$ and $\bm{V}_{R}^{*}$. In the first case and $p=t=3$,
$\bm{V}_{1}^{*}$ $=\bm{V}_{1}^{-1}-\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{V}_{1}^{-1}\bm{{1}}_{p}\right)^{-1}\bm{V}_{1}^{-1}\bm{J}_{p\times p}\bm{V}_{1}^{-1}=\gamma_{11}^{-1}\frac{1}{3-r_{(1)}}\begin{bmatrix}2&-1-r_{(1)}&-1+r_{(1)}\\ -1-r_{(1)}&2+2r_{(1)}&-1-r_{(1)}\\ -1+r_{(1)}&-1-r_{(1)}&2\end{bmatrix}$

### v2

@@ 3.2 Generalized Markov-type covariance structure
$\bm{{C}}_{d^{*}(s2)}$ $=\frac{n}{t-1}\begin{bmatrix}\bm{\mathit{\Lambda}}_{1}&\bm{\mathit{\Lambda}}_{2}\\ \bm{\mathit{\Lambda}}_{2}&\bm{\mathit{\Lambda}}_{4}\end{bmatrix},$ (3.5)
where the matrices $\bm{\mathit{\Lambda}}_{1}$, $\bm{\mathit{\Lambda}}_{2}$ and $\bm{\mathit{\Lambda}}_{4}$ are given as
$\begin{split}\bm{\mathit{\Lambda}}_{1}&=\left(tr\left(\bm{V}_{1}^{*}\right)+\frac{\bar{\rho}^{2}}{\sigma_{12}}tr\left(\bm{V}_{R}^{*}\right)-\frac{\left(tr\left(\bm{V}_{1}^{*}\bm{\psi}\right)\right)^{2}}{tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)}-\frac{\bar{\rho}^{2}}{\sigma_{12}}\frac{\left(tr\left(\bm{V}_{R}^{*}\bm{\psi}\right)\right)^{2}}{tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)}\right)\bm{H}_{t},\\ \bm{\mathit{\Lambda}}_{2}&=-\frac{\bar{\rho}}{\sigma_{12}}\left(tr\left(\bm{V}_{R}^{*}\right)-\frac{\left(tr\left(\bm{V}_{R}^{*}\bm{\psi}\right)\right)^{2}}{tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)}\right)\bm{H}_{t},\text{ and}\\ \bm{\mathit{\Lambda}}_{4}&=\frac{1}{\sigma_{12}}\left(tr\left(\bm{V}_{R}^{*}\right)-\frac{\left(tr\left(\bm{V}_{R}^{*}\bm{\psi}\right)\right)^{2}}{tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)}\right)\bm{H}_{t}.\end{split}$ (3.6)
@@
Since both conditions are not satisfied (see proof of Remark 3.2), we instead focus on efficient designs. In order to check whether $d^{*}$ is an efficient design, we obtain an upper bound of $tr\left(\bm{C}_{d(s2)}\right)$. If $d^{*}$ attains this upper bound then we say that $d^{*}$ is an efficient design. However, obtaining an upper bound for $tr\left(\bm{C}_{d(s2)}\right)$ is computationally challenging, so we instead consider $\bm{\tilde{C}}_{d(s2)}$ (see Lemma S.5), the information matrix for the direct effects with no period effects, and show that $\bm{{C}}_{d(s2)}\leq\bm{\tilde{C}}_{d(s2)}$ with equality if $d$ is uniform on periods (see Lemma S.5 and Lemma S.7).
@@
$\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}=\{d:d\in\Omega_{t,n=\lambda t(t-1),p=t},\text{ }t\geq 3\text{ and }d\text{ is a binary design}\}.$
Also, for $p=t$, a binary design is a design uniform on subjects and thus $\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$ is indeed a class of designs with $p=t$ which are uniform on subjects.
@@
$\left(tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)+\frac{1+\bar{\rho}^{2}}{\sigma_{12}}tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)\right)^{-1}\Bigg{)}$. Here $\bm{V}_{1}^{*}$, $\bm{V}_{R}^{*}$, $\bar{\rho}$ and $\sigma_{12}$ are as defined before.
@@
$\begin{split}tr\left(\bm{\tilde{C}}_{d(s2)(11)}\right)&=tr\left(\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{1}\right)\bm{T}_{d}\right)+tr\left(\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{4}\right)\bm{T}_{d}\right),\\ tr\left(\bm{\tilde{C}}_{d(s2)(12)}\right)&=tr\left(\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{1}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\right)+tr\left(\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{4}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\right),\text{ and}\\ tr\left(\bm{\tilde{C}}_{d(s2)(22)}\right)&=tr\left(\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{1}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\right)+tr\left(\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{4}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\right),\end{split}$ (3.11)
from Lemma S.6. Using Remark S.1, it is observed that $\bm{{1}}_{p}^{{}^{\prime}}\bm{\mathit{\Omega}}_{1}=\bm{{1}}_{p}^{{}^{\prime}}\bm{\mathit{\Omega}}_{2}=\bm{{1}}_{p}^{{}^{\prime}}\bm{\mathit{\Omega}}_{4}=\bm{{0}}_{1\times p}$, and
$\begin{split}tr\left(\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{1}\right)\bm{T}_{d}\right)&=n\times tr\left(\bm{\mathit{\Omega}}_{1}\right),\\ tr\left(\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{4}\right)\bm{T}_{d}\right)&=n\times tr\left(\bm{\mathit{\Omega}}_{4}\right),\\ tr\left(\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{1}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\right)&=n\times tr\left(\bm{\mathit{\Omega}}_{1}\bm{\psi}\right),\\ tr\left(\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{4}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\right)&=n\times tr\left(\bm{\mathit{\Omega}}_{4}\bm{\psi}\right),\\ tr\left(\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{1}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\right)&=n\times tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{1}\bm{\psi}\right),\text{ and}\\ tr\left(\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{4}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\right)&=n\times tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{4}\bm{\psi}\right).\end{split}$ (3.12)
@@
$c_{22(22)}\big{)}$, where $c_{11(11)}=tr\left(\bm{\mathit{\Omega}}_{1}\right)$, $c_{11(22)}=tr\left(\bm{\mathit{\Omega}}_{4}\right)$, $c_{12(11)}=tr\left(\bm{\mathit{\Omega}}_{1}\bm{\psi}\right)$, $c_{12(22)}=tr\left(\bm{\mathit{\Omega}}_{4}\bm{\psi}\right)$, $c_{22(11)}=tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{1}\bm{\psi}\right)$ and $c_{22(22)}=tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{4}\bm{\psi}\right)$. From Lemma S.4, $\bm{\mathit{\Omega}}_{1}=\bm{V}_{1}^{*}+\frac{\bar{\rho}^{2}}{\sigma_{12}}\bm{V}_{R}^{*}$ and $\bm{\mathit{\Omega}}_{4}=\frac{1}{\sigma_{12}}\bm{V}_{R}^{*}$, where $\bar{\rho}=\rho\sqrt{\frac{\sigma_{22}}{\sigma_{11}}}$ and $\sigma_{12}=\sigma_{22}(1-\rho^{2})$. Using these expressions in $c_{22(11)}$ and $c_{22(22)}$, we obtain $c_{22(11)}=tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)+\frac{\bar{\rho}^{2}}{\sigma_{12}}tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)$ and $c_{22(22)}=\frac{1}{\sigma_{12}}tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)$. Using Lemma S.8, $|\bar{\rho}|>0$ and $\sigma_{12}>0$ gives $c_{22(11)}$, $c_{22(22)}>0$, implying $c_{22}=n\left(c_{22(11)}+c_{22(22)}\right)>0$. Combining (3.9), (3.10), (3.11) and (3.12), we obtain
@@
Note that since $c_{22}>0$, $c_{22}^{+}$ has been replaced by $c_{22}^{-1}$. Using the expression of $c_{11}$, $c_{12}$ and $c_{22}$, and the expressions of $\bm{\mathit{\Omega}}_{1}$ and $\bm{\mathit{\Omega}}_{4}$ in (3.13), we prove (3.8). ∎
@@
Let $d^{*}\in\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$ be a design given by $OA_{I}\left(n=\lambda t\left(t-1\right),p=t,t,2\right)$, where $\lambda$ is a positive integer and $t\geq 3$. Then $d^{*}$ attains $u\left(t,n,p,\sigma_{11},\sigma_{22},\rho,\bm{V}_{1},\bm{V}_{R}\right)$ as given in Lemma 3.1, where $n=\lambda t\left(t-1\right)$ and $p=t$, if and only if $tr\left(\bm{V}_{1}^{*}\bm{\psi}\right)tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)=tr\left(\bm{V}_{R}^{*}\bm{\psi}\right)tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)$. Here, $\bm{V}_{1}^{*}$ and $\bm{V}_{R}^{*}$ are as defined before.
@@
The proof of the above remark is given in B. ∎
@@ 3.2.1 Study of design efficiency for the Markovian covariance structure
In this section, for various choices for $\bm{V}_{1}$ and $\bm{V}_{R}$ we evaluate $d^{*}$ using the upper bound of $tr\left(\bm{C}_{d(s2)}\right)$. For the purpose of simplicity, we denote the upper bound given in Lemma 3.1 by $u$. We list $7$ cases based on different combinations of $\bm{V}_{1}$ and $\bm{V}_{R}$ from Table 2:
  1. Case $1$.
    $\bm{V}_{1}$: Mat($0.5$) and $\bm{V}_{R}$: Mat($1.5$)
  2. Case $2$.
    $\bm{V}_{1}$: Mat($0.5$) and $\bm{V}_{R}$: Mat($\infty$)
  3. Case $3$.
    $\bm{V}_{1}$: Mat($1.5$) and $\bm{V}_{R}$: Mat($0.5$)
  4. Case $4$.
    $\bm{V}_{1}$: Mat($1.5$) and $\bm{V}_{R}$: Mat($\infty$)
  1. Case $5$.
    $\bm{V}_{1}$: Mat($\infty$) and $\bm{V}_{R}$: Mat($0.5$)
  2. Case $6$.
    $\bm{V}_{1}$: Mat($\infty$) and $\bm{V}_{R}$: Mat($1.5$)
  3. Case $7$.
    $\bm{V}_{1}$: Mat($0.5$) and $\bm{V}_{R}$: Mat($0.5$)
Note in case $7$, $\bm{\mathit{\Sigma}}$ resembles the NS1 structure, where both $\bm{V}_{1}$ and $\bm{V}_{R}$ have a Mat($0.5$) like structure with correlation parameters $r_{(1)}$ and $r_{(1)}^{2}$; $10<r_{(1)}<$ 1, respectively.
For our illustrations, $p=t\in\{3,4\}$ is used. To apply Remark 3.3, the expressions for $\bm{V}_{1}^{*}$ and $\bm{V}_{R}^{*}$ are needed. In case $7$ for $p=t=3$,
$\bm{V}_{1}^{*}$ $=\bm{V}_{1}^{-1}-\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{V}_{1}^{-1}\bm{{1}}_{p}\right)^{-1}\bm{V}_{1}^{-1}\bm{J}_{p\times p}\bm{V}_{1}^{-1}$
$=\frac{\sigma_{11}^{-1}}{r_{(1)}^{2}-4r_{(1)}+3}\begin{bmatrix}\frac{2\left(r_{(1)}^{2}-4r_{(1)}+3\right)}{r_{(1)}^{3}-3r_{(1)}^{2}-r_{(1)}+3}&-1&\frac{r_{(1)}^{2}-4r_{(1)}+3}{r_{(1)}^{2}-2r_{(1)}-3}\\ -1&2&-1\\ \frac{r_{(1)}^{2}-4r_{(1)}+3}{r_{(1)}^{2}-2r_{(1)}-3}&-1&\frac{2\left(r_{(1)}^{2}-4r_{(1)}+3\right)}{r_{(1)}^{3}-3r_{(1)}^{2}-r_{(1)}+3}\end{bmatrix}$